A 20kg load is dragged along a horizontal surface, if the coefficient of kinetic friction between the load and the surface is 0.23, what is the magnitude of the force applied at angle 30° above the horizontal that results to the load having an acceleration of 0.75m/s² (g= 9.8m/s²)


1
Expert's answer
2021-02-28T07:38:08-0500
"Ox: F-\mu N-mgcos \alpha=ma"

"Oy:N-mgsin \alpha=0"

"F=m(a+gcos\alpha+\mu g sin\alpha)"

"F=20(0.75+9.8\times0.866+0.23\times9.8\times0.5)=207.3 N"